# Water Ice in the Edge-On Orion Silhouette Disk 114–426 from JWST NIRCam Images

### v1

@@ 4.1 Red Edge of the Water Ice Feature
The ice absorption models we present in Section 3.3 result in 3 features that do not extend to long enough wavelengths to reproduce the depth seen in the F335M band (Figure 7). An extended red edge to this feature is also routinely observed in spectra towards protostars and through background clouds (e.g., Smith et al., 1989; Dartois et al., 2002; Noble et al., 2013), including most recently with JWST/NIRSpec (McClure et al., 2023). Dartois et al. (2024) demonstrated how to successfully fit the extended red edge of these NIRSpec spectra. Their model is more sophisticated than ours in three ways. (1) They include additional ice species beyond water, e.g., CH3OH and NH3. (2) They derive the dust optical properties using the discrete dipole approximation to account for more realistic grain shapes. (3) They use the RADMC-3D code (Dullemond et al., 2012) to model the radiation transfer of the background photons through the cloud. Such procedures are beyond the scope of this work and not warranted given the coarse spectral information in the photometric data. However, if future observations of 114–426 at higher spectral resolution detect the extended red edge, these more sophisticated modeling approaches would be justified.
4.2 Origins of the Pa- Absorption
As described in Section 3.2, the absorption spectra in the silhouette of the disk show dips in the F182M and F187N filters. Given that the H II and PDR nebular background is bright in atomic hydrogen lines (including Pa-), and that we are unaware of any plausible dust or ice opacity feature at this wavelength, it seems most likely that the dip in the absorption spectra in these bands is due to preferential absorption of Pa- photons.
Pa- absorption is a transition from the n=3 to n=4 electronic states of atomic hydrogen. Is it plausible that the disk hosts a reservoir of hydrogen in the n=3 state? It is unlikely that gas in the outer disk is hot enough to populate the n=3 level thermally. The gas may be excited from the ground state by absorption of Lyman- photons. These could arise from the central star via its chromosphere or accretion shocks. Alternatively, Lyman- photons could arise from massive stars in the ONC or diffuse nebular radiation. 114–426 does not display the cometary-shaped ionization front surrounding the disk that is typical of photoevaporating “proplyd” disks seen throughout the ONC, meaning this system is shielded from most ionizing EUV photons. However, Miotello et al. (2012) argued that 114–426 does receive a significant dose of external FUV radiation, which includes the Lyman- line.
@@ 4.3 Grain Growth in the Outer Disk
Our finding of grains with radii up to 1 in the outer regions of the disk confirms that dust has grown larger than the sizes found in the ISM. This maximum size is also consistent with ALMA images of 114–426, which found that the sub-mm continuum (tracing even larger grains) only extends throughout the scattered light region of the disk and not into the silhouette (Mann et al., 2014; Bally et al., 2015). The lack of pebbles in the outer regions could be because: (1) grain growth beyond a few microns is inefficient at large disk radii, (2) the large grains that formed in this region have spiralled inwards due to aerodynamic drag (Weidenschilling, 1977), or (3) the outer ansae represent outflowing material in which only small grains are entrained (Miotello et al., 2012).
The grain sizes we derive differ compared to the results of some prior studies of the 114–426 silhouette. Throop et al. (2001) found no difference in absorption between H- (0.656 ) and Pa- images, and so concluded that grains had grown to sizes larger than 5 . Shuping et al. (2003) extended this analysis by detecting somewhat less absorption in a ground-based Br- (4.05 ) image, and concluded the grains were 1.9–4 in size. Our finding of somewhat smaller grains is likely due to our use of continuum bands (and exclusion of the Pa- band) when deriving grain sizes, whereas these prior studies exclusively used images in narrow filters centered on the atomic hydrogen lines to maximize contrast of the silhouettes against the background nebula. As shown in Figures 7 and 8, our models fit a rise in the absorption spectra from 1–2 and ignore the drop in the F187N filter. Models relying on the F187N filter would not fit this rise, inferring a flatter absorption spectrum and thus larger grains. Miotello et al. (2012) fit the absorption in five HST filters, four of which had wide bandwidths, and found 0.2–0.7 grains in the northeast silhouette. Our results agree with those findings while using data out to longer wavelengths.
@@ 4.4 Survivability of Water Ice
The ONC hosts luminous OB stars that bathe the other cluster members and their disks with a strong radiation field. 114–426 resides at a projected separation of 0.2 pc from the massive star Ori C in the Trapezium region, but the line-of-sight separation is unknown so the true separation may be significantly larger. Miotello et al. (2012) estimate that 114–426 experiences an external FUV flux of , although this too is very uncertain. is defined as 108 photons cm-2 s-1 and is approximately equal to the strength of the external FUV field in the local ISM around the sun (Habing, 1968). 114–426 is likely exposed to a level of external irradiation significantly higher than experienced by disks in low-mass star-forming regions (e.g., Taurus, Lupus, Chamaeleon) but lower than experienced by Trapezium Cluster members closer to the center of the Orion Nebula often associated with ionized proplyds (). The key question to answer is whether it is reasonable to find water ice in the outer parts of a disk in this intermediate environment.
@@
One impact of an external radiation field is that it may heat the disk above the sublimation temperature of volatile species, keeping them in the gas phase. Haworth (2021) investigated this possibility and found that the water snow line remains largely immune from external heating if the disk resides more than 0.02 pc from Ori C, which is the case for 114–426. Their models found the midplane temperature in the outer disk of a system like 114–426 to be 50–60 K. This value is below the sublimation temperature of water (150 K), comparable to the sublimation temperature of CO2 (50 K), and above the sublimation temperature of CO (20 K). ALMA detected CO gas from 114–426 in absorption against the background nebula (Bally et al., 2015). The background CO has a brightness temperature of 50–70 K, and the optically-thick CO in the disk was seen to be 12–27 K below this level. Thus, observations of the gas temperature are consistent with the disk being cold enough to maintain water ice against thermal desorption.
@@
More importantly, the external UV field may non-thermally photodesorb water molecules off of the grains and into the gas phase. The efficiency of this process depends on the strength of the FUV field within the icy regions of the disk. To make an order of magnitude estimate, we adopt the incident FUV flux from Miotello et al. (2012) and quantify the extinction arising from the disk material itself, . The factor of 0.5 arises because we are interested in the attenuation of the external field halfway through the total column density. This simplification does not take into account photons entering from varying angles (Cleeves et al., 2013), but it provides a reasonable estimate assuming most photons enter from the direction of the background nebula. Here is the absorption opacity of the dust in the FUV. We compute the opacity spectrum from 912–2000Å with the optool code using the dust properties found when fitting to the NIR data in Section 3.3. We average the opacity spectrum over this wavelength range, weighted by a 39,000 K blackbody (the effective temperature of Ori C) to approximate the spectral shape of the external field in the core of the Orion Nebula. Figure 9 shows the predicted UV field at each position within the silhouette of the disk. The outer less-attenuated (and less icy) radii have spanning 100–1000 . In the inner, denser regions, drops to 10–100 .
@@
Given this FUV field strength, do we expect most of the water to be in ice or gas phase? Motivated by the formalisms presented by Hollenbach et al. (2009) and Öberg et al. (2009), we can estimate the equilibrium amount of water ice versus vapor by equating the photodesorption rate (per volume) of ice into gas with the adsorption rate (per volume) of gas into ice:
@@
There are many significant unknowns regarding the physical and chemical conditions of the 114–426 outer disk, leading to values of the critical ranging from 10 to 10,000 in the parameter space we explore. However, we find that for fiducial values of and , and gas density 108, the majority of the water can remain in the ice even in the outer parts of the disk where the incident FUV field is unattenuated. Thus, we conclude that it is entirely plausible for water ice to survive against the effects of UV photodesorption in this environment, and the detection of water ice is not at odds with what we know about this disk.
@@
The ability of water ice to survive in externally irradiated disks like 114–426 is significant because most stars, including the Sun, form in cluster environments like the ONC (Lada & Lada, 2003; Winter & Haworth, 2022), and an external FUV flux of 102–104 is the most typical field strength experienced by young stars (Fatuzzo & Adams, 2008). Two other edge-on disks further from the center of the Orion Nebula (132–1832 and 216–0939) have 3 water ice detections (Terada & Tokunaga, 2012; Terada et al., 2012). With the addition of water ice detected in 114–426, there is growing evidence for water ice in intermediately-irradiated regions of clustered star formation – the typical environment of planet formation in our Galaxy.
@@ 5 Summary
We present an analysis of the 114–426 edge-on silhouette disk in 12 JWST NIRCam images. 114–426 is large on the sky and resides in front of the background emission of the Orion Nebula. This offers a unique opportunity to study the disk material in absorption. The NIRCam images span the prominent 3 water ice absorption feature, allowing us to search for this important volatile in a reservoir of planet-forming material. Our main findings include:
@@
We fit a model absorption spectrum of water ice-coated grains to each pixel in the silhouette. The ice-to-refractory mass ratio ranges from 0 to 0.18, the maximum grain size ranges from 0.25 to 5 , and the total mass of solids in this region is 0.46 . Observations at higher spectral resolution would improve the accuracy of these findings.
We also observe a dip in the absorption spectra at the wavelength of Pa-, suggesting there is excited atomic hydrogen in the absorbing material. The gas may be excited by UV photons from the central star or external sources, but a quantitative assessment will require future study.
@@
This study provides direct observational evidence that water ice can survive in an intermediately-irradiated disk, the typical planet-forming environment in the Galaxy.
Higher spectral resolution near-IR observations of this unique system would confirm many of the results presented here and open the door to further discoveries. The NIRSpec instrument onboard JWST, for instance, has an IFU mode that would be ideally suited for this task. At higher resolution the water ice absorption feature could be better isolated from the neighboring continuum and its spectral profile measured, enabling a more accurate determination of the ice abundance. The shape of the profile can indicate whether the water ice is amorphous or crystalline (Mastrapa et al., 2009). An extended red edge to the feature, hinted at in the NIRCam data, can also reveal if the ice is pure or mixed with other species such as CH3OH and NH3 (Dartois et al., 2024). Furthermore, such observations could search for the distinct narrower absorption features from more volatile CO2 and CO ice at 4.27 and 4.67 . Finally, higher spectral resolution observations would be ideal to detect absorption (or emission) from gas phase atomic and molecular species in the silhouette, notably the Pa- line seen in this study.
We thank the referee for their feedback that improved the clarify of the paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data used are available in the Barbara A. Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The dataset ID is http://dx.doi.org/10.17909/vjys-x251.
@@ Appendix A Results after image convolution
Figure 11 shows the gallery of images convolved to the F470N resolution. Differences are minimal with respect to the corresponding unconvolved images. The wavelength-dependent brightness asymmetry and lateral offset between the two scattered light lobes remains apparent. The shrinking of the dark midplane with increasing wavelength is also still evident, although the trend is somewhat counteracted by the higher degree of smoothing in the shorter wavelength images.
The spectra extracted at certain pixels from the convolved images are shown in gray in Figures 4 and 5. They reveal approximately the same pattern as the unconvolved spectra (shown in black) – including the Pa- and water ice features – except they are shifted to slightly higher or lower values. All of the absorption spectra in the silhouette are shifted up, indicating less absorption. This is expected, as more flux from the surrounding nebula is blurred into the silhouette. The spectra from the scattered light lobes (U, V, X, Y) are fainter after convolution because some of the flux is blurred away. The spectra from the midplane (W, Z) are slightly brighter because additional flux from the scattered light lobes is blurred into this region.
We fit the absorption spectra of all pixels in the silhouette of the convolved images using the same method as for the unconvolved images (Section 3.3), and the maps of the best-fit parameters are shown in Figure 12. We find the ice-to-refractory mass ratio is generally higher by up to 50%, the maximum grain size agrees to within 40%, and the solid surface density is 10–30% lower when using the convolved images compared to the unconvolved images. Overall, we conclude that the variation in spatial resolution across the NIRCam bands used for this analysis does not significantly bias the results presented in the main text.

### v2

@@ 4.1 Red Edge of the Water Ice Feature
The ice absorption models we present in Section 3.3 result in 3 features that do not extend to long enough wavelengths to reproduce the depth seen in the F335M band (Figure 7). An extended red edge to this feature is also routinely observed in spectra toward protostars and through background clouds (e.g., Smith et al., 1989; Dartois et al., 2002; Noble et al., 2013), including most recently with JWST/NIRSpec (McClure et al., 2023). Dartois et al. (2024) demonstrated how to successfully fit the extended red edge of these NIRSpec spectra. Their model is more sophisticated than ours in three ways: (1) they include additional ice species beyond water, e.g., CH3OH and NH3, (2) they derive the dust optical properties using the discrete dipole approximation to account for more realistic grain shapes, (3) they use the RADMC-3D code (Dullemond et al., 2012) to model the radiation transfer of the background photons through the cloud. Such procedures are beyond the scope of this work and not warranted given the coarse spectral information in the photometric data. However, if future observations of 114–426 at higher spectral resolution detect the extended red edge, these more sophisticated modeling approaches would be justified.
4.2 Origins of the Pa Absorption
As described in Section 3.2, the absorption spectra in the silhouette of the disk show dips in the F182M and F187N filters. Given that the H II and PDR nebular background is bright in atomic hydrogen lines (including Pa), and that we are unaware of any plausible dust or ice opacity feature at this wavelength, it seems most likely that the dip in the absorption spectra in these bands is due to preferential absorption of Pa photons.
Pa absorption is a transition from the =3 to =4 electronic states of atomic hydrogen. Is it plausible that the disk hosts a reservoir of hydrogen in the =3 state? It is unlikely that gas in the outer disk is hot enough to populate the =3 level thermally. The gas may be excited from the ground state by the absorption of Ly photons. These could arise from the central star via its chromosphere or accretion shocks. Alternatively, Ly photons could arise from massive stars in the ONC or diffuse nebular radiation. 114–426 does not display the cometary-shaped ionization front surrounding the disk that is typical of photoevaporating “proplyd” disks seen throughout the ONC, meaning this system is shielded from most ionizing extreme-ultraviolet photons. However, Miotello et al. (2012) argued that 114–426 does receive a significant dose of external far-ultraviolet (FUV) radiation, which includes the Ly line.
@@ 4.3 Grain Growth in the Outer Disk
Our finding of grains with radii up to 1 in the outer regions of the disk confirms that dust has grown larger than the sizes found in the interstellar medium (ISM). This maximum size is also consistent with ALMA images of 114–426, which found that the submillimeter continuum (tracing even larger grains) only extends throughout the scattered light region of the disk and not into the silhouette (Mann et al., 2014; Bally et al., 2015). The lack of pebbles in the outer regions could be because (1) grain growth beyond a few microns is inefficient at large disk radii, (2) the large grains that formed in this region have spiraled inward due to aerodynamic drag (Weidenschilling, 1977), or (3) the outer ansae represent outflowing material in which only small grains are entrained (Miotello et al., 2012).
The grain sizes we derive differ compared to the results of some prior studies of the 114–426 silhouette. Throop et al. (2001) found no difference in absorption between H (0.656 ) and Pa images, and so concluded that grains had grown to sizes larger than 5 . Shuping et al. (2003) extended this analysis by detecting somewhat less absorption in a ground-based Br (4.05 ) image, and concluded the grains were 1.9–4 in size. Our finding of somewhat smaller grains is likely due to our use of continuum bands (and exclusion of the Pa band) when deriving grain sizes, whereas these prior studies exclusively used images in narrow filters centered on the atomic hydrogen lines to maximize contrast of the silhouettes against the background nebula. As shown in Figures 7 and 8, our models fit a rise in the absorption spectra from 1 to 2 and ignore the drop in the F187N filter. Models relying on the F187N filter would not fit this rise, inferring a flatter absorption spectrum and thus larger grains. Miotello et al. (2012) fit the absorption in five HST filters, four of which had wide bandwidths, and found 0.2–0.7 grains in the northeast silhouette. Our results agree with those findings while using data out to longer wavelengths.
@@ 4.4 Survivability of Water Ice
The ONC hosts luminous OB stars that bathe the other cluster members and their disks with a strong radiation field. 114–426 resides at a projected separation of 0.2 pc from the massive star Ori C in the Trapezium region, but the line-of-sight separation is unknown so the true separation may be significantly larger. Miotello et al. (2012) estimate that 114–426 experiences an external FUV flux of , although this too is very uncertain. is defined as 108 photons cm-2 s-1 and is approximately equal to the strength of the external FUV field in the local ISM around the sun (Habing, 1968). 114–426 is likely exposed to a level of external irradiation significantly higher than experienced by disks in low-mass star-forming regions (e.g., Taurus, Lupus, and Chamaeleon) but lower than experienced by Trapezium Cluster members closer to the center of the Orion Nebula often associated with ionized proplyds (). The key question to answer is whether it is reasonable to find water ice in the outer parts of a disk in this intermediate environment.
@@
One impact of an external radiation field is that it may heat the disk above the sublimation temperature of volatile species, keeping them in the gas phase. Haworth (2021) investigated this possibility and found that the water snow line remains largely immune from external heating if the disk resides more than 0.02 pc from Ori C, which is the case for 114–426. Their models found the midplane temperature in the outer disk of a system like 114–426 to be 50–60 K. This value is below the sublimation temperature of water (150 K), comparable to the sublimation temperature of CO2 (50 K), and above the sublimation temperature of CO (20 K). ALMA detected CO gas from 114–426 in absorption against the background nebula (Bally et al., 2015). The background CO has a brightness temperature of 50–70 K, and the optically thick CO in the disk was seen to be 12–27 K below this level. Thus, observations of the gas temperature are consistent with the disk being cold enough to maintain water ice against thermal desorption.
@@
More importantly, the external UV field may nonthermally photodesorb water molecules off of the grains and into the gas phase. The efficiency of this process depends on the strength of the FUV field within the icy regions of the disk. To make an order of magnitude estimate, we adopt the incident FUV flux from Miotello et al. (2012) and quantify the extinction arising from the disk material itself, . The factor of 0.5 arises because we are interested in the attenuation of the external field halfway through the total column density. This simplification does not take into account photons entering from varying angles (Cleeves et al., 2013), but it provides a reasonable estimate assuming most photons enter from the direction of the background nebula. Here is the absorption opacity of the dust in the FUV. We compute the opacity spectrum from 912 to 2000Å with the optool code using the dust properties found when fitting to the near-IR data in Section 3.3. We average the opacity spectrum over this wavelength range, weighted by a 39,000 K blackbody (the effective temperature of Ori C) to approximate the spectral shape of the external field in the core of the Orion Nebula. Figure 9 shows the predicted UV field at each position within the silhouette of the disk. The outer less-attenuated (and less icy) radii have spanning 100–1000 . In the inner, denser regions, drops to 10–100 .
@@
Given this FUV field strength, do we expect most of the water to be in ice or gas phase? Motivated by the formalisms presented by Hollenbach et al. (2009) and Öberg et al. (2009), we can estimate the equilibrium amount of water ice versus vapor by equating the photodesorption rate (per volume) of ice into gas with the adsorption rate (per volume) of gas into ice
@@
There are many significant unknowns regarding the physical and chemical conditions of the 114–426 outer disk, leading to values of the critical ranging from 10 to 10,000 in the parameter space we explore. However, we find that for fiducial values of and , and gas density 108 cm-3, the majority of the water can remain as ice even in the outer parts of the disk where the incident FUV field is unattenuated. Thus, we conclude that it is entirely plausible for water ice to survive against the effects of UV photodesorption in this environment, and the detection of water ice is not at odds with what we know about this disk.
@@
The ability of water ice to survive in externally irradiated disks like 114–426 is significant because most stars, including the Sun, form in cluster environments like the ONC (Lada & Lada, 2003; Winter & Haworth, 2022), and an external FUV flux of 102–104 is the most typical field strength experienced by young stars (Fatuzzo & Adams, 2008). Two other edge-on disks further from the center of the Orion Nebula (132–1832 and 216–0939) have 3 water ice detections (Terada & Tokunaga, 2012; Terada et al., 2012). With the addition of water ice detected in 114–426, there is growing evidence for water ice in intermediately irradiated regions of clustered star formation – the typical environment of planet formation in our Galaxy.
@@ 5 Summary
We present an analysis of the 114–426 edge-on silhouette disk in 12 JWST NIRCam images. 114–426 is large on the sky and resides in front of the background emission of the Orion Nebula. This offers a unique opportunity to study the disk material in absorption. The NIRCam images span the prominent 3 water ice absorption feature, allowing us to search for this important volatile in a reservoir of planet-forming material. Our main findings include the following.
@@
We fit a model absorption spectrum of water-ice-coated grains to each pixel in the silhouette. The ice-to-refractory mass ratio ranges from 0 to 0.18, the maximum grain size ranges from 0.25 to 5 , and the total mass of solids in this region is 0.46 . Observations at higher spectral resolution would improve the accuracy of these findings.
We also observe a dip in the absorption spectra at the wavelength of Pa, suggesting there is excited atomic hydrogen in the absorbing material. The gas may be excited by UV photons from the central star or external sources, but a quantitative assessment will require future study.
@@
This study provides direct observational evidence that water ice can survive in an intermediately irradiated disk, the typical planet-forming environment in the Galaxy.
Higher spectral resolution near-IR observations of this unique system would confirm many of the results presented here and open the door to further discoveries. The NIRSpec instrument on board JWST, for instance, has an integral field unit mode that would be ideally suited for this task. At higher resolution the water ice absorption feature could be better isolated from the neighboring continuum and its spectral profile measured, enabling a more accurate determination of the ice abundance. The shape of the profile can indicate whether the water ice is amorphous or crystalline (Mastrapa et al., 2009). An extended red edge to the feature, hinted at in the NIRCam data, can also reveal if the ice is pure or mixed with other species such as CH3OH and NH3 (Dartois et al., 2024). Furthermore, such observations could search for the distinct narrower absorption features from more volatile CO2 and CO ice at 4.27 and 4.67 . Finally, higher spectral resolution observations would be ideal to detect absorption (or emission) from gas phase atomic and molecular species in the silhouette, notably the Pa line seen in this study.
We thank the referee for their feedback that improved the clarity of the paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data used are available in the Barbara A. Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The dataset ID is 10.17909/vjys-x251.
@@ Appendix A Results after image convolution
Figure 11 shows a gallery of images convolved to the F470N resolution. Differences are minimal with respect to the corresponding unconvolved images. The wavelength-dependent brightness asymmetry and lateral offset between the two scattered light lobes remain apparent. The shrinking of the dark midplane with increasing wavelength is also still evident, although the trend is somewhat counteracted by the higher degree of smoothing in the shorter-wavelength images.
The spectra extracted at certain pixels from the convolved images are shown in gray in Figures 4 and 5. They reveal approximately the same pattern as the unconvolved spectra (shown in black) – including the Pa and water ice features – except they are shifted to slightly higher or lower values. All of the absorption spectra in the silhouette are shifted up, indicating less absorption. This is expected, as more flux from the surrounding nebula is blurred into the silhouette. The spectra from the scattered light lobes (U, V, X, and Y) are fainter after convolution because some of the flux is blurred away. The spectra from the midplane (W and Z) are slightly brighter because additional flux from the scattered light lobes is blurred into this region.
We fit the absorption spectra of all pixels in the silhouette of the convolved images using the same method as for the unconvolved images (Section 3.3), and maps of the best-fit parameters are shown in Figure 12. We find the ice-to-refractory mass ratio is generally higher by up to 50%, the maximum grain size agrees to within 40%, and the solid surface density is 10%–30% lower when using the convolved images compared to the unconvolved images. Overall, we conclude that the variation in spatial resolution across the NIRCam bands used for this analysis does not significantly bias the results presented in the main text.